OpenStudy (anonymous):

Suppose the drug atenolol is taken in 25 mg doses once a day to lower blood pressure. At the end of a 24 hour period, approximately 7 % of the atenolol is still present. (a) Find expressions for the quantities Q1, Q2, Q3, . . ., and Qn , representing the amounts of atenolol in the body right after taking the 1st, 2nd, 3rd doses respectively. Then write the expression for Qn in closed-form.

OpenStudy (anonymous):

i think this is like summing a geometric sequence

OpenStudy (anonymous):

first term is 25 second term is \(25+.07\times 25\)

OpenStudy (anonymous):

does this look familiar?

OpenStudy (anonymous):

third term is \[25+.07(25+.07\times 25)=25+.07\times 25+.07^2\times 25\]

OpenStudy (anonymous):

and fourth , cutting to the chase, is \[25+.07\times 25+.07^2\times 25+.07^2\times 25\] or \[25(1+.07+.07^2+.07^3)\]